# PhysLLM: Harnessing Large Language Models for Cross-Modal Remote Physiological Sensing

### v1

@@ 3.2 Physiological Cue-Aware Prompt Learning
$\Gamma(x_{\mathrm{enc}}[b,:]),\ \mathrm{sgn}\left(\sum_{t=1}^{T-1}(x_{\mathrm{enc}}[b,t+1]-x_{\mathrm{enc}}[b,t])\right),$
$\mathrm{TopK}\left(\mathcal{L}(x_{\mathrm{enc}}[b,:]),5\right)\Biggr{\}}$
where $\Gamma(\cdot)$ calculates signal trends through first-order differencing,
$\small\Gamma(\mathbf{x})=\sum_{i=2}^{T}(\mathbf{x}_{i}-\mathbf{x}_{i-1}).\vspace{-0.2em}$ (17)
Then $\mathcal{S}$ will make up the symbolic representation as described in Fig. 4(c). Finally, the tokenization process is formulated as $\mathcal{C}_{static}=\mathrm{Tokenizer}\left(\mathcal{S}\right).$
Adaptive Prompt Learning (APL). Conventional multimodal fusion approaches typically employ static weighting coefficients for prompt integration, suffering from two limitations: (1) fixed combination ratios cannot adapt to learnable LLM, and (2) inability to adaptively select the key cues across different data. As shown in Fig. 2(a), our adaptive fusion paradigm addresses these constraints through learnable hierarchical composition. Formally, having obtained three fundamental cue captions $\mathcal{C}=\{\mathcal{C}_{task},\mathcal{C}_{vision},\mathcal{C}_{static}\}$, we first implement modality-specific conditioning via dedicated transformation networks,
[…]
$\small\operatorname{AttentiveCompressor}_{k}(x)=\operatorname{Softmax}\left(\frac{Q_{k}K_{k}^{T}}{\sqrt{d}}\right)V_{k},\vspace{-0.3em}$ (19)
with ${Q_{k},K_{k},V_{k}}$ derived from $x$ through linear projections.
$\small\mathbf{W}=\left[W_{\text{task}},\ W_{\text{vision}},\ W_{\text{static}}\right]\in\mathbb{R}^{3\times B\times L\times d},\vspace{-0.3em}$ (20)
where $L$ is predefined hyperparameter about target prompt sequence length, $d$ is the token embedding dimension, and $W_{\text{task}},W_{\text{vision}},W_{\text{static}}$ denote learnable tokens, respectively.
The complete fusion operation is formalized as:
$\small\mathcal{T}_{\text{cue }}=\sum_{\begin{subarray}{c}k\in\Omega\end{subarray}}\mathbf{W}^{(k)}\odot\mathcal{E}_{k},\Omega\triangleq\{task,\\ vision,\\ static\}.\vspace{-0.3em}$ (21)
3.3 Training Objectives
In Section 3.1, we obtain the stationary signal $z$ and the fused multi-scale visual features $F_{visual}$ through the DDS module and multi-scale interaction, respectively. Through TPG, we can obtain the text-guided vision token $\mathcal{T}_{\text{vision }}=TPG(F_{visual})$, and the text-guided signal token $\mathcal{T}_{\text{signal }}=TPG(z)$. Then, we input $\mathcal{T}_{\text{cue }}$, $\mathcal{T}_{\text{vision }}$, $\mathcal{T}_{\text{signal }}$ into LLM and predict waveforms $\hat{y}$. Finally, we employ the mean square error loss function as the total loss,
$\small\mathcal{L}_{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\vspace{-0.3em}$ (22)
@@ 4.1 Datasets and Performance Metrics
We comprehensively assess the performance of PhysLLM on four widely used benchmark datasets (UBFC-rPPG [1], PURE [27], BUAA [37], and MMPD [31]), in comparison with state-of-the-art methods. Following [28, 38], we calculate mean absolute error (MAE), root mean square error (RMSE), and Pearson’s correlation coefficient (R) between the predicted HRs versus the ground-truth HRs as evaluation metrics. Notably, for MAE and RMSE, lower values indicate reduced error margins, whereas for R, values approaching 1.0 signify diminished error. Among them, both MAE and RMSE are measured in terms of bpm (beats per minute). And for convenience, these units will not be repeated in the subsequent tables or analyses.
4.2 Implementation Details
We conduct experiments on Pytorch and mainly based on the open-source toolkit rPPG-Toolbox [15]. For data pre-processing, we crop the face region in the first frame for each video clip and fix the region box in the following frames. Subsequently, we randomly sample a video chunk of 128 frames and resize them into 128 × 128 pixels. We use the default hyperparameters settings $\alpha$ = 0.8 and $l_{target}$ = 32. The backbone of our PhysLLM is based on the pre-trained PhysNet [38], where the training strategies follow the methodology outlined in the original paper. In the intra-dataset testing and ablation study, we use the same training data for PhysNet as for PhysLLM, ensuring consistency in the data used. For cross-domain experiments, we use the pre-trained PhysNet model trained on the PURE dataset [27], ensuring no test set overlap and consistent pre-training conditions. Furthermore, we use the Deepseek-1.5B [7] version as the LLM. The PhysLLM is trained with Adam optimizer and the initial learning rate and weight decay are 1e-4 and 5e-5, respectively. We train our model for 20 epochs on a NVIDIA A100 GPU with batch size of 4.
@@ 4.2 Implementation Details
Efficientphys [13] 11.9 18.5 11.8 18.9 15.5 20.8
PhysFormer [41] 13.9 18.6 11.4 17.5 13.2 16.5
PhysNet [39] 13.2 16.7 11.0 17.3 13.5 17.0
@@
[…]
Efficientphys [13] 13.2 20.02 0.08 32.3 34.0 -0.03
PhysFormer [41] 13.9 19.3 0.06 7.65 12.4 0.25
PhysNet [38] 12.8 16.3 0.12 12.75 16.37 0.08
PhysLLM (Ours) 12.2 15.5 0.19 6.0 8.6 0.63
4.3 Intra-dataset Testing
We first evaluate the HR estimation on all datasets under intra-dataset setting. We compare our method with 10 methods, including traditional methods, Green [33], ICA [22], CHROM [5], POS [35], and deep learning methods, Meta-rPPG [9], Contrast-Phys+ [29], PhysNet [39], PhysFormer [41], Efficientphys [13]. Table 1 displays intra-dataset testing results for UBFC-rPPG [1], PURE [27], BUAA [37] and MMPD [31].
HR Estimation on UBFC-rPPG [1]. On the UBFC-rPPG [1] dataset, we use the videos of the first 30 subjects for training and the videos of the remaining 12 subjects for testing according to  [18]. The HR estimation results are shown in Table 1. The proposed PhysLLM outperforms the state-of-the-art methods, and achieves MAE (0.21 bpm), RMSE (0.57 bpm), and R (0.99) metrics for HR prediction. The results show that PhysLLM not only fits the rPPG signal consistently but also tracks HR changes effectively over a long period.
HR Estimation on PURE [27]. Following the protocol in  [18], we compare PhysLLM with 9 methods. It can be seen from Table 1 that PhysLLM achieves the best performance on all evaluation metrics for HR, especially compared with the second-best method PhysFormer [41]. PhysLLM achieves improvements of 0.31 bpm in MAE and 0.63 bpm in RMSE, indicating the excellent robustness of PhysLLM against the interference induced by head movements.
HR Estimation on BUAA [37]. It is partitioned in sequence into training and test sets with a ratio of 7:3. The performance of existing methods on this dataset is reimplemented by ourselves with the rPPG toolbox [15]. The HR estimation results are shown in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (6.48 bpm), RMSE (8.48 bpm), and R (0.63) metrics for HR prediction. The results show the strong robustness of PhysLLM against various lighting condition disturbances.
HR Estimation on MMPD [31]. Following the protocol in  [46], we compared our method against existing state-of-the-art approaches, as depicted in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (4.36 bpm), RMSE (10.76 bpm), and R (0.65) metrics for HR prediction. These results demonstrate that PhysLLM performs well under real-world conditions.
@@ 4.4 Cross-dataset Testing
We also perform cross-dataset testing across UBFC-rPPG [1], PURE [27], BUAA [37], and MMPD [31] datasets to assess the generalization capability of PhysLLM. We conduct both Two-training and One-testing protocol and Three-training and One-testing protocol. In each case, we train on relatively simple datasets and evaluate performance on one more complex dataset.
Table 4: Ablation results of the main components on UBFC-rPPG [1].
Method PURE
@@
Dual-source training and single-target testing protocol.  To explore the generalization capacity of PhysLLM, we train on two datasets and uniformly test on the most difficult dataset, MMPD [31]. For example, $P+U\to M$ means training on PURE [27] and BUAA [37] datasets and testing on MMPD [31], and both the training and testing take the entire dataset. The two-training and one-testing dataset HR estimation results of our method and baseline methods are presented in Table 2. Despite more difficult cross-domain settings, PhysLLM still achieves the best performance, even reducing the MAE by 1.05 bpm and the RMSE by 2.34 bpm on $P+U\to M$, reflecting the superior generalization capacity of PhysLLM even on complex unseen domains, and the effectiveness of the visual contexts and LLM modeling.
Three-source training and single-target testing protocol. To further evaluate whether PhysLLM can learn domain invariance knowledge without mixing domain-specific knowledge, we trained on three datasets and tested on the remaining one dataset. For example, $Others\to MMPD$ means training on the other datasets (UBFC-rPPG [1], PURE [27], BUAA [37]) and test on MMPD [31]. It is worth noting that we only tested two difficult datasets: BUAA [37] and MMPD [31]. It can be seen from the results in Table 3 that PhysLLM achieves the best performance in both datasets, and even outperforms the existing state-of-the-art methods on MAE (6.0 bpm), RMSE (8.6 bpm), and R (0.63) metrics for HR prediction on $Others\to BUAA$. This shows that PhysLLM will not confuse domain-specific knowledge even if trained on three datasets, and can stably learn domain-invariant rPPG knowledge.
Table 5: Ablation study of prompt components on UBFC-rPPG [1].
@@
4.5 Ablation Study
Impact of Text Prototype Guidance (TPG). To evaluate the effectiveness of text prototype guidance, we conducted an ablation study by removing the text prototype guidance. The results on UBFC-rPPG dataset are shown in the fourth row of Table  4. TPG brings a significant decrease in MAE (0.06 bpm) and RMSE (0.35 bpm). These results underscore the importance of SFD in facilitating the fusion of quasi-periodic video features.
Impact of prompt components. To evaluate the influence of different prompt components (Vision, Statics, Task, and adaptive learning) in cue caption and adaptive prompt learning of PhysLLM, we conducted ablation on them and presented in Table. 5 using MAE and RMSE metrics. The results reveal that the optimal configuration includes all the components achieving remarkable performance with MAE 0.21 bpm and RMSE 0.57 bpm. This represents a substantial 76% improvement over the worst-performing combination (MAE 1.31 bpm, RMSE 2.10 bpm), which utilized only Vision, Task and Statics components while excluding adaptive learning. This shows that our adaptive learning part is essential and critical. Several significant patterns emerge: configurations incorporating the Task component generally outperform those without it, suggesting the critical importance of task-specific prompting; the combination of Vision and Task is higher than that of Statics and Task, suggesting the generation of visual prompt is more in line with the relevant rules of rPPG tasks. Moderate performance (0.71-0.84 MAE) is achieved with Statics and adaptive learning only, but none approach the effectiveness of the optimal Task and adaptive learning pairing. These findings demonstrate the crucial importance of careful prompt component selection and emphasize that task-specific adaptability should be prioritized in prompt design, while also emphasizing the importance of learning headers assigning different prompt weights to different datasets and scenarios.
Figure 5: Visualization of saliency maps from PhysLLM on PURE [27] and UBFC-rPPG [1].
@@ 4.6 Visualization and Discussion
Following [28], the saliency maps of the PhysLLM are visualized in Fig. 5 using samples from UBFC-rPPG [1] and PURE [27] datasets. The more obvious the blue-green area, the more attention it means. We selected images with certain features from the dataset to test rPPG tasks such as head rotation, with a noticeable beard, darker skin tone, and wearing glasses. From the test results, it can be seen that physLLM can effectively capture rPPG-related parts. For example, the observable parts are concentrated on the cheeks and forehead, which is consistent with the relevant prior knowledge of rPPG. At the same time, relevant parts can effectively avoid hair and obstruction. For example, there is no prominent attention in the area where the hair appears, and even subtle changes in the neck can be observed. This is enough to reflect the robustness and effectiveness of PhysLLM in different scenarios.
5 Conclusions and Future Work

### v2

@@ 3.3 Training Objectives
$\mathcal{L}_{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2},\vskip-10.00002pt$ (20)
@@ 4.1 Datasets and Performance Metrics
We comprehensively evaluate models on four benchmark datasets (UBFC-rPPG (Bobbia et al., 2019), PURE (Stricker et al., 2014), BUAA (Xi et al., 2020), and MMPD (Tang et al., 2023)). Following (Sun and Li, 2022; Yu et al., 2019), we calculate mean absolute error (MAE), root mean square error (RMSE), and Pearson’s correlation coefficient (R) between the predicted HRs versus the ground-truth HRs as evaluation metrics. Notably, for MAE and RMSE, lower values indicate reduced error margins, whereas for R, values approaching 1.0 signify diminished error. Among them, both MAE and RMSE are measured in terms of bpm (beats per minute). More implementation details can be found in Appendix. B.
4.2 Intra-dataset Testing
We first evaluate the HR estimation on all datasets under intra-dataset setting. We compare our method with 10 methods, including traditional methods and deep learning methods. Table 1 displays intra-dataset testing results for UBFC-rPPG (Bobbia et al., 2019), PURE (Stricker et al., 2014), BUAA (Xi et al., 2020) and MMPD (Tang et al., 2023).
HR Estimation on UBFC-rPPG (Bobbia et al., 2019). On UBFC-rPPG (Bobbia et al., 2019), we follow  (Luo et al., 2024) by training on the first 30 subjects and testing on the remaining 12. As shown in Table 1, PhysLLM achieves state-of-the-art HR estimation with MAE 0.21 bpm, RMSE 0.57 bpm, and R 0.99. These results indicate stable rPPG fitting and effective long-term HR tracking.
HR Estimation on PURE (Stricker et al., 2014). Following  (Luo et al., 2024), we compare PhysLLM with 9 methods. As shown in Table 1, PhysLLM achieves the best HR performance across all metrics, outperforming the second-best PhysFormer (Yu et al., 2022) by 0.31 bpm MAE and 0.63 bpm RMSE, demonstrating strong robustness to head-motion interference.
HR Estimation on BUAA (Xi et al., 2020). It is partitioned in sequence into training and test sets with a ratio of 7:3. The performance of existing methods on this dataset is reimplemented by ourselves with the rPPG toolbox (Liu et al., 2023c). The HR estimation results are shown in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (6.48 bpm), RMSE (8.48 bpm), and R (0.63) metrics for HR prediction. The results show the strong robustness of PhysLLM against various lighting condition disturbances.
@@ 4.2 Intra-dataset Testing
Green (Verkruysse et al., 2008) 21.7 27.7 21.7 27.7 21.7 27.7
EfficientPhys (Liu et al., 2021a) 11.9 18.5 11.8 18.9 15.5 20.8
PhysFormer (Yu et al., 2022) 13.9 18.6 11.4 17.5 13.2 16.5
PhysNet (Yu et al., 2019) 13.2 16.7 11.0 17.3 13.5 17.0
RhythmFormer (Zou et al., 2025) 13.98 19.5 10.5 16.7 12.6 17.5
@@
$downarrow MAE\downarrow RMSE\downarrow$
Green (Verkruysse et al., 2008) 21.7 27.7 6.9 10.4
EfficientPhys (Liu et al., 2021a) 13.2 20.02 32.3 34.0
PhysFormer (Yu et al., 2022) 13.9 19.3 7.7 12.4
PhysNet (Yu et al., 2019) 12.8 16.3 12.8 16.4
RhythmFormer (Zou et al., 2025) 16.1 20.5 6.04 10.8
PhysLLM (Ours) 12.2 15.5 6.01 8.6
HR Estimation on MMPD (Tang et al., 2023). Following the protocol in  (Zou et al., 2024), we compared our method against existing state-of-the-art approaches, as depicted in Table 1. The proposed PhysLLM outperforms the existing state-of-the-art methods on MAE (4.36 bpm), RMSE (10.76 bpm), and R (0.65) metrics for HR prediction. These results demonstrate that PhysLLM performs well under real-world conditions.
@@ 4.3 Cross-dataset Testing
We also perform cross-dataset testing to assess the generalization capability of PhysLLM. We conduct both Two-training and One-testing protocol and Three-training and One-testing protocol. In each case, we train on relatively simple datasets and evaluate performance on one more complex dataset.
Dual-source training and single-target testing protocol.  To evaluate generalization, we train on two datasets and test on the most challenging MMPD (Tang et al., 2023). For instance, $P+U\to M$ denotes training on PURE (Stricker et al., 2014) and BUAA (Xi et al., 2020) and testing on MMPD using the full datasets. As shown in Table 3, PhysLLM remains best under these cross-domain settings, reducing MAE/RMSE by 1.05/2.34 bpm on $P+U\to M$, demonstrating strong generalization enabled by visual context and LLM modeling.
Table 4: Ablation results of the main components on the UBFC-rPPG (Bobbia et al., 2019) dataset.
Method UBFC-rPPG
@@
Three-source training and single-target testing protocol. To further evaluate whether PhysLLM can learn domain invariance knowledge without mixing domain-specific knowledge, we trained on three datasets and tested on the remaining one dataset. For example, $Others\to MMPD$ means training on the other datasets (UBFC-rPPG (Bobbia et al., 2019), PURE (Stricker et al., 2014), BUAA (Xi et al., 2020)) and test on MMPD (Tang et al., 2023). It is worth noting that we only tested two difficult datasets: BUAA (Xi et al., 2020) and MMPD (Tang et al., 2023). It can be seen from the results in Table 3 that PhysLLM achieves the best performance in both datasets, and even outperforms the existing state-of-the-art methods on MAE (6.0 bpm), RMSE (8.6 bpm), and R (0.63) metrics for HR prediction on $Others\to BUAA$. This shows that PhysLLM will not confuse domain-specific knowledge even if trained on three datasets, and can stably learn domain-invariant rPPG knowledge.
Table 5: Ablation experiments of different large language models and transformer baseline. While DeepSeek is used as the default LLM in our experiments, the proposed approach can be easily generalized to other large language models.
Method LLM Param. UBFC PURE
$downarrow MAE\downarrow RMSE\downarrow$
PhysNet (Yu et al., 2019) - 2.95 3.67 2.10 2.60
PhysLLM (w. Sundial (Liu et al., 2025)) - 0.92 2.46 3.22 7.54
PhysLLM (w. DeepSeek (Guo et al., 2025)) 1.5B 0.21 0.57 0.17 0.35
PhysLLM (w. Bert (Devlin et al., 2019)) 0.11B 0.19 0.76 0.43 0.80
PhysLLM (w. GPT2 (Lagler et al., 2013)) 0.124B 0.19 0.76 0.14 0.35
4.4 Ablation Study
Impact of Text Prototype Guidance (TPG). To evaluate the effectiveness of text prototype guidance, we conducted an ablation study by removing the text prototype guidance. The results on UBFC-rPPG dataset are shown in the fourth row of Table 4. TPG brings a significant decrease in MAE (0.06 bpm) and RMSE (0.35 bpm). These results underscore the importance of TPG in facilitating the fusion of quasi-periodic video features.
Table 6: Ablation study of prompt components on the UBFC-rPPG (Bobbia et al., 2019) dataset.
@@ 4.4 Ablation Study
Impact of prompt components. To evaluate the contributions of different prompt components (Vision, Statics, Task, and adaptive learning) in PhysLLM’s cue caption and adaptive prompt learning, we conduct an ablation study and report MAE/RMSE in Table 6. The full configuration achieves the best performance (MAE 0.21 bpm, RMSE 0.57 bpm), yielding a 76% improvement over the worst setting (MAE 1.31 bpm, RMSE 2.10 bpm) that removes adaptive learning, confirming its critical role. We also observe that adding the Task component generally improves performance, and Vision+Task outperforms Statics+Task, suggesting visual prompts better align with rPPG-specific cues.
Ablation Study on the LLM Component. To validate the necessity of using a pre-trained LLM, we compare three LLMs with varying sizes: DeepSeek (Guo et al., 2025), Bert (Devlin et al., 2019), and GPT2 (Lagler et al., 2013). As shown in Table 5, all LLMs achieve competitive performance, with DeepSeek performing best (MAE: 0.21/0.17 on UBFC/PURE). We further replace the LLM with Sundial (Liu et al., 2025), a temporal transformer without LLM pre-training. The significant performance drop (MAE: 3.35/4.05) demonstrates that the LLM’s pre-trained knowledge is essential for cross-dataset generalization, not merely the transformer architecture.
@@ 4.5 Visualization and Discussion
Robustness Analysis on Skin Tones and Lighting Conditions. To evaluate robustness, we perform stress tests on the MMPD (Tang et al., 2023) dataset by varying skin tones (Types 3–6) and lighting conditions (LED-Low/High, Incandescent, Natural). As shown in Table 8, PhysLLM consistently outperforms PhysFormer and RhythmFormer across all settings, especially under extreme lighting (MAE/RMSE: 3.45/6.81 in natural light) and diverse skin tones (4.73/8.49 for Type 4). Overall, PhysLLM remains stable under challenging variations, supporting its real-world applicability across different populations and environments.
Model Complexity Analysis. As shown in Table 8, PhysLLM has higher computational complexity (97.2M parameters, 424.3G MACs) due to the LLM backbone. While we acknowledge this overhead, it is a necessary trade-off for achieving superior cross-dataset generalization and robustness demonstrated in our experiments. Future work will explore model compression techniques such as knowledge distillation and parameter-efficient fine-tuning to reduce costs while maintaining performance for resource-constrained deployment.
Figure 5: Visualization of saliency maps from PhysLLM on PURE and UBFC-rPPG datasets.
@@
Visualization of saliency maps. Following (Sun and Li, 2022), the saliency maps of the PhysLLM are visualized in Fig. 5 using samples from UBFC-rPPG (Bobbia et al., 2019) and PURE (Stricker et al., 2014) datasets. The more obvious the red-green area, the more attention it means. We selected images with certain features from the dataset to test rPPG tasks, such as head rotation, with a noticeable beard, darker skin tone, and wearing glasses. From the test results, it can be seen that PhysLLM can effectively capture rPPG-related parts. For example, the observable parts are concentrated on the cheeks and forehead, which is consistent with the relevant prior knowledge of rPPG. At the same time, relevant parts can effectively avoid hair and obstruction. For example, there is no prominent attention in the area where the hair appears, and even subtle changes in the neck can be observed. This is sufficient to demonstrate the robustness and effectiveness of PhysLLM in various scenarios.
Table 7: Performance Comparison across Different Skin Tones and Lighting Conditions on MMPD (Tang et al., 2023) dataset (MAE/RMSE).
(a) Performance across Different Skin Tones.
Method Type 3 Type 4 Type 5 Type 6
PhysFormer (Yu et al., 2022) 6.11/11.21 5.92/10.00 5.87/13.12 6.81/11.21
RhythmFormer (Zou et al., 2025) 5.12/11.23 5.46/9.14 6.32/11.87 6.26/9.99
PhysLLM (Ours) 4.96/10.27 4.73/8.49 4.99/11.07 5.73/8.49
(b) Performance across Different Lighting Conditions.
Method LED-Low LED-High Incandescent Natural
PhysFormer (Yu et al., 2022) 6.36/11.72 5.12/9.39 5.71/12.73 6.61/11.36
RhythmFormer (Zou et al., 2025) 5.85/11.71 4.46/8.98 3.64/11.87 5.65/12.31
PhysLLM (Ours) 4.46/9.57 3.72/8.10 3.59/11.43 3.45/6.81
Table 8: Comparison of Model Complexity
Method Params (M) MACs (G)
TS-CAN (Liu et al., 2020) 7.5 96.0
PhysNet (Yu et al., 2019) 0.77 56.1
DeepPhys (Chen and McDuff, 2018) 7.5 96.0
EfficientPhys (Liu et al., 2021a) 7.4 45.6
PhysFormer (Yu et al., 2022) 7.38 40.5
RhythmFormer (Zou et al., 2025) 4.21 28.8
Contrast-phys+ (Sun and Li, 2024) 0.85 145.7
PhysMamba (Luo et al., 2024) 0.56 47.3
PhysLLM (Ours) 97.2 424.3
5 Conclusion